OpenStudy (anonymous):

HELP PLEASE! MEDAL GIVEN! In a population of grasshoppers, the color green (G) is completely dominant over the color brown (g). If 23% of the population is brown, calculate the percentage of the population that is expected to be heterozygous (Gg). 50 percent 47 percent 77 percent 35 percent

OpenStudy (anonymous):

I don't think its 77 percent as http://openstudy.com/study#/updates/508b1975e4b077c2ef2e7671 says otherwise

OpenStudy (anonymous):

but i don't know the answer

OpenStudy (anonymous):

f (gg) =q ^2 = 0.23 so f(g) = q= 0.48 p= 1-q = 0.52 Frequency of heterozygotes is 2pq = 0.5 = 50%

OpenStudy (anonymous):

OK, thank you so much!

OpenStudy (anonymous):

Welcome!